OpenStudy (anonymous):

Use the quadratic formula to find the x-intercepts of the function. Round to the nearest tenth if necessary. y = -2x2 + 5x - 1 A. {2.3, 0.2} B. {2.2, -5.8} C. {-9.2, -0.8} D. no real solution

OpenStudy (anonymous):

\[\frac{5\pm\sqrt{17}}{4}\]

OpenStudy (anonymous):

best answer is A

OpenStudy (anonymous):

Hmmm alright (: Thanks!!!

OpenStudy (anonymous):

Can you help me on like my last 6 questions?

OpenStudy (anonymous):

i just computed first by quadratic formula, then plug in to calculator

OpenStudy (anonymous):

ok shoot

OpenStudy (anonymous):

pleaseeeeee? u dont have to.. but i would like it!

OpenStudy (anonymous):

i am ready, lets do it.

OpenStudy (anonymous):

Solve, using the square root method: 5x2 = 60. Round the solution to the nearest tenth. A. {3.5, -3.5} B. {7.7, -7.7} C. {3.5} D. No real solution

OpenStudy (anonymous):

(: So generous!! :D

OpenStudy (anonymous):

3.5, -3.5

OpenStudy (anonymous):

5 to go

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Okay next one Solve using the square root method. Round to the nearest tenth if necessary. x2 + 15 =0 A. {3.9} B. {3.9, -3.9} C. {7.5, -7.5} D. no real solution

OpenStudy (anonymous):

no real solution since \[x^2=-15\] has no real answer

OpenStudy (anonymous):

4 to go

OpenStudy (anonymous):

OKay so D Next one Which number completes the perfect square trinomial? x2 + 20x + ? A. 40 B. 100 C. 25 D. 80

OpenStudy (anonymous):

100

OpenStudy (anonymous):

(:

OpenStudy (anonymous):

half of 20 is 10, 10 squared is 100

OpenStudy (anonymous):

3 more

OpenStudy (anonymous):

Okay (: Use the quadratic formula to solve the problem. A soccer player throws the ball onto the field from the sideline. An equation modeling the ball's flight is h = -16t2 + 18t + 2. In how many seconds will the ball reach the ground? A. in about 2 seconds B. in about 1.2 seconds C. in about 1.8 seconds D. Cannot be determined

OpenStudy (anonymous):

set \[-16t^2+18t+2=0\] solve for t

OpenStudy (anonymous):

B

OpenStudy (anonymous):

about 1.2

OpenStudy (anonymous):

Haha yea alright (: Thanks soooo much btw (: Next.. Which of the following shows the square completed in the this equation? x2 + 3x - 5 = 0 A. (x + 9)2 = 14 B. (x - 3)2 = 8 C. (x + 1.5)2 = 7.25

OpenStudy (anonymous):

this is nonsense by the way unless he throws it straight up in the air. if he throws it out on the field equation makes no sense

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Yea... thats true!!

OpenStudy (anonymous):

one more?

OpenStudy (anonymous):

2 more! here they are Solve the following using the quadratic formula. Round to the nearest tenth if necessary. 2x2 + 9x + 4 = 0 A. {-4, -0.5} B. {-16, -2} C. {-8, -1} D. no real solution Use the quadratic formula to find the zeros of the function y = 2x2 + 3x + 6. Round to the nearest tenth if necessary. A. {-1.7, 2.3} B. {-2.4, 2.4} C. {-2.3, 0.8} D. no real solution

OpenStudy (anonymous):

ohh u dont have to do the last one then if u dont want to (:

OpenStudy (anonymous):

i miscounted my bad! ^^

OpenStudy (anonymous):

first one is A, -4 and -.5

OpenStudy (anonymous):

no problem we do the last one now

OpenStudy (anonymous):

Okay (:

OpenStudy (anonymous):

ahww thanks!! =3

OpenStudy (anonymous):

no solution

OpenStudy (anonymous):

OKay (: OMFG i could seriously stalk you right now!!! (lmfao) I got another 100% I am totally getting it back to a B!!!! And its all thanks to youuuuuuuuu!!! <3

OpenStudy (anonymous):

are you taking this on line?

OpenStudy (anonymous):

Yep!! (; Haha

OpenStudy (anonymous):

way to go!

OpenStudy (anonymous):

Its hard as hell!!!

OpenStudy (anonymous):

how can you miss? just ask and you get the answers.

OpenStudy (anonymous):

and now i am done for the evening. gnight!

OpenStudy (anonymous):

Ikr!! haha i usually get those ppl who only want metals :P And they give me wrong answerssss..... and okay thanks sooooooooooooooooooooo much!!!!!! (: Hope i see u again!! (:

OpenStudy (anonymous):

ok maybe tomorrow.

OpenStudy (anonymous):

very welcome.

OpenStudy (anonymous):

(:

OpenStudy (anonymous):

:)